Sorry, you need to enable JavaScript to visit this website.

MCST Banner Image

am1weekly_2000-008_2000-015

Mission Operations Days: 2000/008 to 2000/015
January 08, 2000 00:00:00 GMT to January 15, 2000 00:00:00 GMT


Terra Spacecraft and MODIS Instrument Status:

Terra (AM-1) is in Normal Ops Mode
Terra (AM-1) has an anomaly with the Propulsion Thrusters
MODIS is in Standby Mode
MODIS has no known Anomalies
ComponentState(s)Comment
BlackbodyA Off; B OffNominal
Calibration ElectronicsA On; B OffNominal
Control ProcessorA Off; B OnNominal
Nadir DoorUnlatched, ClosedNominal
Space View DoorUnlatched, At OutgasNominal
Solar Diffuser DoorUnlatched, ClosedNominal
FDDI FormatterA Off; B OffNominal
FDDI PortA Off; B OffNominal
FIFO Memory#1 & #2 Off; #3 & #4 OffNominal
Format ProcessorA Off; B OffNominal
Power Supplies#1 Off, #2 OnNominal
PV VIS FPAA Off; B OffNominal
PV NIR FPAA Off; B OffNominal
PV SMIR FPAA Off; B OffNominal
PV LWIR FPAA Off; B OffNominal
PC LWIR FPAA Off; B OffNominal
RC Outgas HeatersAll On (CS, IS, OS)Nominal
RC LWIR FPA HeaterOffNominal
RC SMIR FPA HeaterOffNominal
Scan AssemblyA Off; B OffNominal
SDSMOffNominal
SRCAOffNominal
PS1 Survival HeaterEnabledNominal
PS2 Survival HeaterEnabledNominal
Timing GeneratorA Off; B OffNominal
Flight SoftwareRev BDNominal
Inhibit Ids SetNoneNominal
TMONs EnabledNoneNominal

This Week’s Completed MODIS Activities:

Sunday, January 9, 2000

None

Monday, January 10, 2000

None

Tuesday, January 11, 2000

2000/011 23:20 Autonomously sent to Safe Mode by S/C

Wednesday, January 12, 2000

2000/012 21:14 Transitioned to Standby Mode
2000/012 21:17 Turned on the cold and intermediate stage outgas heater
2000/012 21:24 Turned on the outer stage outgas heater

Thursday, January 13, 2000

None

Friday, January 14, 2000

None

Saturday, January 15, 2000

None

This Week's Scheduled MODIS Activities Not Completed:

None

Upcoming MODIS Events:

2000/017 12:30 EST Turn off outgas heaters

Wednesday, January 19, the results of the SBRS thermal model will be reviewed. Based on this discussion a decision will be made when to transition to Science Mode.

MODIS Anomalies:

MODIS had a yellow limit, MOD_VR_PS2_P15V_A2AF, at 2000:015:04:57:02 for one cycle. The value was 17.1959V and the yellow upper limit violated is 17.15V. The upper red limit for this point is 18.75V. Plots were made and faxed to JAM who agreed this was worth watching but no action should be taken.

The formatter engine temperatures (A and B) continues to cyclicly drop below their yellow limits. There are no health and safety risks. These temperatures are varying cyclically with the MEM survival heater duty cycle.

General Instrument Comments:

MODIS is currently in Standby Mode. All three MODIS doors have successfully been unlatched. The MEM survival heater cycle is steadily on approximately 40% of the time. MODIS began outgassing on January 3rd, 2000, stopped on January 11, and resumed approximately 22 hours later on January 12.

MODIS Telemetry Trends:

The outgas heater stage temperatures are oscillating as expected under thermostatic control. The 88V power supply feed is fluctuating, generally between 82 and 90 Volts, as each of the three outgas heaters power on and turn off. Similar fluctuations are seen for the S/C power system current feed to MODIS as the outgas heaters toggle. The A and B side formatter engine temperatures are oscillating in conjunction with the MEM survival heater duty cycle and have been exceeding their yellow low limits on a predictable basis.

Non-MODIS Significant Events:

2000/009: ASTER successfully activated their Thermal Infrared Radiometer (TIR) today (one of three separate instruments comprising the ASTER package).

2000/010: Today was the first Staging Burn (also referred to as a Calibration Burn) for the Space Craft. The burn took place at 17:55:42 GMT and lasted 11 seconds.

2000/011: The CPHTS (cooling system) for TIR had a problem and TIR had to be shutdown. The problem is referred to as depriming.

2000/011: S/C autonomously sent to Safe Mode during orbit attainment thruster burn at 18:20EST. Anomaly team continues to work on a recovery plan.

2000/013: Communication with the S/C via the HGA resumed.

2000/013: The TIR CPHTS was restarted on the redundant side. The ASTER team will wait until they are comfortable that the CPHTS is working before turning on TIR again. (CPHTS = Capillary Pump Heat Transfer System)

2000/015: The TIR CPHTS deprimed again.

Limited Life Item Status:

No limited life items were exercised this week. All limited life items are well within lifetime ranges, although precise statistics for each item are still pending.